OpenStudy (richyw):

Volume by DOUBLE integral. Still struggling with the same double integral problem! Test is in two hours! Can someone check my attempt?

OpenStudy (richyw):

set up the double integral to evaluate the volume of the solid bounded by \(y=x^2\), \(z=3y\) and \(z=2+y\)

OpenStudy (richyw):

I have set it up like this. does it look correct? \[V=\int^1_0\int^{2+y}_{3y}x^2\,dzdx\]

OpenStudy (amistre64):

how did you determine the solid that is bounded?

OpenStudy (richyw):

I plugged \(y=x^2\) into \(z=3y\) and \(z=2+y\). And found \(x=\pm 1\).

OpenStudy (richyw):

I am very confused, all of the text book questions I breezed through and this one has me second guessing...

OpenStudy (amistre64):